# Bibliography.bib

### v1

@@ 2.3. Triangulations
We now recall the notion of a triangulation of a manifold; i.e. a decomposition of the manifold into smooth simplices. This combinatorial description of the manifold is particularly useful for local arguments and, as we shall see, will turn our arguments into arguments about linear polyhedra. In this section we fix our notation and conventions on the topic. For a more detailed account of triangulations we refer to [whitney2012geometric, Ch. IV.B] and [lurie2009topics].
@@ 2.3.1. Polyhedra
It is a result of Whitehead [Whitehead1940] that any smooth manifold can be triangulated, uniquely up to subdivision. The name “Whitehead triangulation” is sometimes used to emphasize the compatibility between the triangulation and the smooth structure of $M$.
@@ 2.3.2. Piecewise smooth sections
We denote the space of such sections by $\Gamma_{\operatorname{PS}}(E)$. As before, if we want to emphasize the role of $T$, we may say that $(s,T)$ is a piecewise smooth section.
To define the $C^{r}$-topology on $\Gamma_{\operatorname{PS}}(E)$, we again need to fix a metric on the jet space $J^{r}(E)$. Compared to the setting of maps over linear polyhedra from Definition 2.12, we however do not need to fix a triangulation to make sense of jet bundles over $M$. Then we obtain:
@@
Given a pair of sections $(s_{1},T_{1}),(s_{2},T_{2})\in\Gamma_{\operatorname{PS}}(E)$ we write $d_{C^{r}}(s_{1},s_{2})<\varepsilon$ if there exists a common subdivision $T:|K|\to M$ of $T_{1}$ and $T_{2}$ such that for every simplex $\Delta\in T$ we have $d_{C^{r}}(s_{1}|_{\Delta},s_{2}|_{\Delta})<\varepsilon$. Here we let $\varepsilon$ be a positive constant if $P$ is compact, and a positive function $P\to{\mathbb{R}}_{+}$ if $P$ is not compact.
We point out that, even though the distances defined above will generally differ from those defined in Definition 2.12, the induced topologies on $\Gamma_{\operatorname{PS}}(E)$ do coincide with those in Definition 2.12.
@@ 2.4. Simplicial sets
In this section we briefly recall the necessary theory of simplicial sets. We emphasize that this is a very short overview which does not give a general account of the theory, but only covers the minimum to reach what we need. In particular, we work our way up to the specific characterization of weak homotopy equivalences of simplicial sets appearing in Definition 2.25. This section is mostly based on [heuts2022simplicial, Ch. 2], to which we also refer for more details.
We start by defining the simplex category $\bm{\Delta}$ which has as objects $[n]=\{0,1,\dots,n\}$. Using the natural ordering $0\leq 1\leq\dots\leq n$ we define morphisms in $\bm{\Delta}$ as non-decreasing functions. These are generated by the face maps $\delta^{i}:[n-1]\to[n]$ and the degeneracy maps $\sigma^{i}:[n+1]\to[n]$, which are defined by either skipping or doubling the element $i$, respectively. This allows us to define:
@@
[…]
@@ 2.4.1. Weak homotopy equivalences
In the category of topological spaces a weak homotopy equivalence can be defined as the vanishing of the relative homotopy groups. We can use a similar definition for simplicial sets, but for this we need simplicial analogues of the disc and its boundary. These roles are played by the standard $n$-simplex and its boundary. We use the notation $\Delta[n]$ for the standard $n$-simplex as a simplicial set, and the notation $\Delta^{n}$ as a topological space.
@@
$(\partial\Delta[n])_{m}=\{f\in\operatorname{Hom}_{\bm{\Delta}}([m],[n])\mid f\text{ is not surjective}\}.$
We have two observations to make concerning the above definition. First, note that $\Delta[n]$ and $\operatorname{Sing}(\Delta^{n})$ are not the same. One could show for instance that the latter is a so-called Kan complex whereas the former is not. The geometric realization of $\Delta[n]$ is however the standard simplex $\Delta^{n}$, which explains the notation. Secondly, we note that the Yoneda lemma gives us a bijection $\operatorname{Hom}_{\textbf{sSet}}(\Delta[n],X)\to X_{n}$ for all $X\in{\textbf{sSet}}$. This provides a useful interpretation of morphisms with domain $\Delta[n]$ and clarifies its role in the theory.
@@
The definition of weak homotopy equivalence for simplicial sets stated below is not the standard one but equivalent to it. It is based on a characterization of Jardine [jardine1987simplical]. It generalizes the notion for spaces via the singular complex and the geometric realization functor.
@@
A morphism $f:X\to Y$ of simplicial sets is a weak homotopy equivalence if and only if for all $n\in\mathbb{N}$ and all commutative diagrams
${\partial\Delta[n]}$${X}$${\Delta[n]}$${Y,}$$\scriptstyle{f}$$\scriptstyle{\sigma}$
@@ 3.1. Crystalline subdivision
There are various ways of defining crystalline subdivision, but here we follow the definition of Thurston from [Th2, p. 227].
@@
Figure 1. The second crystalline subdivision of the standard $2$-simplex $\Delta^{2}$.
@@
Applying crystalline subdivision to the standard $m$-simplex $\Delta^{m}$ with $m>0$ is done using the following steps:
@@
    Include $\Delta^{m}$ into the standard $m$-cube $I^{m}$.
@@
This procedure makes use of the ordering of the vertices in step 1. Indeed, for $m\geq 3$, the resulting subdivision would be different if we applied an affine isomorphism that permutes the vertices.
This procedure can be iterated but it turns out to be easier to define finer crystalline subdivisions directly. Hence we define the $\ell$th crystalline subdivision of the standard $m$-simplex $\Delta^{m}$ by instead subdividing the $m$-cube into $2^{\ell m}$ smaller $m$-cubes of size $2^{-\ell}$ in item 2 from Definition 3.1.
@@
Let $K$ be an ordered simplicial complex. Its $\ell$th crystalline subdivision $K_{\ell}$ is obtained as follows: Given a simplex $\Delta\in K$, consider the unique order-preserving, affine identification $\phi$ between $\Delta$ and the standard simplex $\Delta^{m}$. Use $\phi$ to transfer the $\ell$th crystalline subdivision of $\Delta^{m}$ to $\Delta$.
Given a polyhedron $T:|K|\rightarrow M$ associated to a specified simplicial complex $K$, we can also speak of its $\ell$th crystalline subdivision $T_{\ell}$, seen as the collection of embeddings $(T_{\Delta}:\Delta\to M)_{\Delta\in K_{\ell}}$ given by restricting $T$.
@@ 3.2. Properties of crystalline subdivision
We discuss some of the properties of crystalline subdivision next. They all have the same flavour: since crystalline does not distort simplices, we are able to obtain various quantitative bounds independent of the order of subdivision $\ell$.
@@ 3.2.1. Coloring
Consider the standard $m$-cube $I^{m}$ subdivided into smaller $m$-cubes of length $2^{-\ell}$ for some $\ell\in\mathbb{N}$. Let $Q\subset I^{m}$ be one of those smaller cubes. Then $Q$ is adjacent to at most $13^{m}-$ 1 cubes. In the procedure of crystalline subdivision each of these cubes is subdivided into $m!$ smaller $m$-simplices. If $A$ is the number of simplices in $K^{(\mathrm{top})}$, then each simplex $\Delta\in K^{(\mathrm{top})}_{\ell}$ is adjacent to at most $(3^{m}-1)\cdot(m)!\cdot A$ simplices, which means that there exists a $((3^{m}-1)\cdot(m)!\cdot A)$-coloring of $K_{\ell}$. ∎
@@ 3.2.2. Model simplices
A second property of crystalline subdivision is that the resulting simplices are equivalent to the standard simplex, up to scaling and translation. Model simplices appear already in Thurston’s work [Th2], albeit in a different formulation.
@@
Let $K$ be a finite and ordered simplicial complex in ${\mathbb{R}}^{N}$ of pure dimension. Then, there exists a finite collection of model simplices ${\mathcal{C}}=\{\Delta_{i}\subset{\mathbb{R}}^{N}\mid i=0,\dots,I\}$ with the following property: for any $\ell\in\mathbb{N}$ and $\Delta\in K_{\ell}^{(\mathrm{top})}$ there exists
@@
Let $\Delta\in K^{(\mathrm{top})}$ be a top dimensional simplex, then there exists an affine isomorphism $\phi:\Delta\to\Delta^{m}\subset{\mathbb{R}}^{m}$, where $m$ is the dimension of $\Delta$. We define ${\mathcal{C}}^{\prime}_{\Delta}$ to be the first crystalline subdivision of $\Delta^{m}$. We note that any simplex in $(\Delta^{m})_{\ell}$ is by construction a translated and scaled copy of a unique simplex in $(\Delta^{m})_{1}$. Using $\phi$ we now push forward ${\mathcal{C}}^{\prime}_{\Delta}$ to a collection of model simplices ${\mathcal{C}}_{\Delta}$. We define ${\mathcal{C}}$ to be the union over all ${\mathcal{C}}_{\Delta}$ for $\Delta\in K^{(\mathrm{top})}$. ∎
@@ 3.2.3. Shape of simplices
Given a simplex, we are interested in the maximal ($\operatorname{r_{\mathrm{max}}}$) and minimal ($\operatorname{r_{\mathrm{min}}}$) distances between its vertices. The former quantity controls for instance how well a map is approximated by its linearization with respect to a given triangulation (see Section 4). Their ratio on the other hand tells us how degenerate the simplex is. We define these quantities for $(m+1)$-tuples of points in ${\mathbb{R}}^{N}$, but we can also speak of $\operatorname{r_{\mathrm{max}}}$ and $\operatorname{r_{\mathrm{min}}}$ of a linear simplex $\Delta$ by identifying $\Delta$ with its vertices.
@@
The continuous functions $\operatorname{r_{\mathrm{min}}},\operatorname{r_{\mathrm{max}}}:{\mathbb{R}}^{N}\times\dots\times{\mathbb{R}}^{N}={\mathbb{R}}^{N(m+1)}\to{\mathbb{R}}$ are defined by
$\operatorname{r_{\mathrm{min}}}(p_{0},\dots,p_{m})$ $=\min_{i,j\in[m]}d(p_{i},p_{j})\text{ \quad and}$
$\operatorname{r_{\mathrm{max}}}(p_{0},\dots,p_{m})$ $=\max_{i,j\in[m]}d(p_{i},p_{j}).$
A priori we cannot bound these quantities and their ratio among all subdivisions of a linear polyhedron, since the space of all linear simplices in ${\mathbb{R}}^{N}$ is not compact. However, due to the existence of the model simplices (from Lemma 3.4), we can bound these quantities when considering crystalline subdivisions:
@@
Let $K$ be simplicial complex that is ordered and finite. Then, there exists $C,D\in{\mathbb{R}}_{+}$ such that for all $\ell\in\mathbb{N}$:
$\max_{\Delta\in K_{\ell}}\operatorname{r_{\mathrm{max}}}(\Delta)=C\cdot 2^{-\ell}\quad\textrm{ and }\quad\dfrac{\max_{\Delta\in K_{\ell}}\operatorname{r_{\mathrm{max}}}(\Delta)}{\min_{\Delta\in K_{\ell}}\operatorname{r_{\mathrm{min}}}(\Delta)}=D.$

### v2

@@ 2.3. Triangulations
We now recall the notion of a triangulation of a manifold; i.e. a decomposition of the manifold into smooth simplices. This combinatorial description of the manifold is particularly useful for local arguments and, as we shall see, will turn our arguments into arguments about linear polyhedra. In this section we fix our notation and conventions on the topic. For a more detailed account of triangulations we refer to [34, Ch. IV.B] and [21].
@@ 2.3.1. Polyhedra
It is a result of Whitehead [33] that any smooth manifold can be triangulated, which is unique up to a piecewise linear homeomorphism. The name “Whitehead triangulation” is sometimes used to emphasize the compatibility between the triangulation and the smooth structure of $M$.
@@ 2.3.2. Piecewise smooth sections
We denote the space of such sections by $\Gamma_{\mathrm{PS}}(E)$. As before, if we want to emphasize the role of $T$, we may say that $(s,T)$ is a piecewise smooth section.
To define the $C^{r}$-topology on $\Gamma_{\mathrm{PS}}(E)$, we again need to fix a metric on the jet space $J^{r}(E)$. Compared to the setting of maps over linear polyhedra from Definition 2.12, we however do not need to fix a triangulation to make sense of jet bundles over $M$. Then we obtain:
@@
Given a pair of sections $(s_{1},T_{1}),(s_{2},T_{2})\in\Gamma_{\mathrm{PS}}(E)$ we write $d_{C^{r}}(s_{1},s_{2})<\varepsilon$ if there exists a common subdivision $T:|K|\to M$ of $T_{1}$ and $T_{2}$ such that for every simplex $\Delta\in T$ we have $d_{C^{r}}(s_{1}|_{\Delta},s_{2}|_{\Delta})<\varepsilon$. Here we let $\varepsilon$ be a positive constant if $P$ is compact, and a positive function $P\to{\mathbb{R}}_{+}$ if $P$ is not compact.
We point out that, even though the distances defined above will generally differ from those defined in Definition 2.12, the induced topologies on $\Gamma_{\mathrm{PS}}(E)$ do coincide with those in Definition 2.12.
@@ 2.4. Simplicial sets
In this section we briefly recall the necessary theory of simplicial sets. We emphasize that this is a very short overview which does not give a general account of the theory, but only covers the minimum to reach what we need. In particular, we work our way up to the specific characterization of weak homotopy equivalences of simplicial sets appearing in Definition 2.29. This section is mostly based on [16, Ch. 2], to which we also refer for more details.
We start by defining the simplex category $\bm{\Delta}$ which has objects $n]=\{0,1,\dots,n\}$. Using the natural ordering $0\leq 1\leq\dots\leq n$ we define morphisms in $\bm{\Delta}$ as non-decreasing functions. These are generated by the face maps $\delta^{i}:[n-1]\to[n]$ and the degeneracy maps $\sigma^{i}:[n+1]\to[n]$, which are defined by either skipping or doubling the element $i$, respectively. This allows us to define:
@@
The geometric realization $|\cdot|:{\textbf{sSet}}\to{\textbf{Top}}$ functor is also relevant to our discussion. We do not want to define it precisely, but we do want to point out that it can be seen as the inverse of the singular complex in two ways. Firstly, it holds that there exists a bijection
@@ 2.4.1. Weak homotopy equivalences
In the category of topological spaces a weak homotopy equivalence can be defined as the vanishing of the relative homotopy groups. We can use a similar definition for (a subcategory of) simplicial sets, but for this we need simplicial analogues of the disc and its boundary. These roles are played by the standard $n$-simplex and its boundary. We use the notation $\Delta[n]$ for the standard $n$-simplex as a simplicial set, and the notation $\Delta^{n}$ as a topological space.
@@
[…]
We have two observations to make concerning the above definition. First, note that $\Delta[n]$ and $\operatorname{Sing}(\Delta^{n})$ are not the same. One could show for instance that the latter is a Kan complex, which we introduce below, whereas the former is not. The geometric realization of $\Delta[n]$ is however the standard simplex $\Delta^{n}$, which explains the notation. Secondly, we note that the Yoneda lemma gives us a bijection $\operatorname{Hom}_{\textbf{sSet}}(\Delta[n],X)\to X_{n}$ for all $X\in{\textbf{sSet}}$. This provides a useful interpretation of morphisms with domain $\Delta[n]$ and clarifies its role in the theory.
@@
To define weak homotopy equivalences for simplicial sets, we restrict ourselves here to Kan complexes, which are simplicial sets that allow for so-called horn-fillings. As a space, we define the $k$-horn $\Lambda^{n}_{k}$ of $\Delta^{n}$ as the subcomplex $\overline{(\partial\Delta^{n}\setminus F_{k})}$ of the standard $n$-simplex $\Delta$ where $F_{k}$ is the face of $\Delta^{n}$ excluding the $k$th vertex. It is the geometric realization of its counterpart in simplicial sets, which we denote by $\Lambda[n,k]$.
@@
The $k$-horn of the standard $n$-simplex $\Lambda[n,k]$ is defined as a simplicial set by
$(\Lambda[n,k])_{i}=\{f\in\operatorname{Hom}_{\Delta}([i],[n])\mid{\operatorname{Im}}f\cup\{k\}\neq[n]\}.$
An example of a Kan complex to keep in mind is the singular complex of a topological space.
Definition 2.28.
A simplicial set $X$ is a Kan complex if for every $n\geq 1$ and $k\in[n]$ any morphism $\Lambda[n,k]\to X$ can be extended to a morphism $\Delta[n]\to X$.
Restricting ourselves to Kan complexes allows us to define weak homotopy equivalences for simplicial sets, in a way that resembles the case for topological spaces. The definition stated below is not the standard one but equivalent to it in the case of Kan complexes [18, 10]. It generalizes the notion for spaces via the singular complex and the geometric realization functor.
Definition 2.29.
A morphism $f:X\to Y$ of Kan complexes is a weak homotopy equivalence if and only if for all $n\in\mathbb{N}$ and all commutative diagrams
[…]
@@ 3.1. Crystalline subdivision
There are various ways of defining crystalline subdivision, but here we follow the definition of Thurston from [31, p. 227].
@@
Figure 1. The second crystalline subdivision of a $2$-simplex $\Delta$.
@@
Applying crystalline subdivision to an (ordered) linear $m$-simplex $\Delta$ with $m>0$ is done using the following steps:
@@
    Include $\Delta=\langle v_{i_{0}},\dots,v_{i_{m}}\rangle$ into the standard $m$-cube $I^{m}$ via the piecewise linear map $\iota$ defined by
    $\iota(v_{i_{j}})=({\overbrace{0,\dots,0}^{j}},{\overbrace{1,\dots,1}^{m-j}}).$
@@
  4. (4)
    Use $\iota$ to pullback the above subdivision of $\iota(\Delta)$ to $\Delta$.
@@
This procedure makes use of the ordering of the vertices in step 1. Indeed, for $m\geq 3$, the resulting subdivision would be different if we applied an affine isomorphism that permutes the vertices of the simplex $\Delta$. It is possible to assign an ordering to the result of crystalline subdivision of $\Delta$, but as we have no use for this, we do not.
[…]
@@
[…]
@@ 3.2. Properties of crystalline subdivision
We discuss some of the properties of crystalline subdivision next. They all have the same flavor: since crystalline does not distort simplices, we are able to obtain various quantitative bounds independent of the order of subdivision $\ell$.
@@ 3.2.1. Coloring
Consider the standard $m$-cube $I^{m}$ subdivided into smaller $m$-cubes of length $2^{-\ell}$ for some $\ell\in\mathbb{N}$. Let $Q\subset I^{m}$ be one of those smaller cubes. Then $Q$ is adjacent to at most $3^{m}-1$ cubes. In the procedure of crystalline subdivision each of these cubes is subdivided into $m!$ smaller $m$-simplices. If $A$ is the number of simplices in $K^{(\mathrm{top})}$, then each simplex $\Delta\in K^{(\mathrm{top})}_{\ell}$ is adjacent to at most $(3^{m}-1)\cdot(m)!\cdot A$ simplices, which means that there exists a $((3^{m}-1)\cdot(m)!\cdot A)$-coloring of $K_{\ell}$. ∎
@@ 3.2.2. Model simplices
[…]
@@
Let $K$ be a finite, ordered simplicial complex of pure dimension $m$. Then, there exists a finite collection of model simplices ${\mathcal{C}}=\{\Delta_{i}\subset{\mathbb{R}}^{N}\mid i=0,\dots,I\}$ with the following property: for any $\ell\in\mathbb{N}$ and $\Delta\in K_{\ell}^{(\mathrm{top})}$ there exists
@@
Let $\Delta\in K^{(\mathrm{top})}$ be a top-dimensional simplex, then we define ${\mathcal{C}}^{\prime}_{\Delta}$ to be the first crystalline subdivision of $\Delta$. We note that any simplex in $(\Delta)_{\ell}$ is by construction a translated and scaled copy of a unique simplex in $(\Delta)_{1}$. We define $\mathcal{C}$} to be the union over all ${\mathcal{C}}_{\Delta}$ for $\Delta\in K^{(\mathrm{top})}$. ∎
@@ 3.2.3. Shape of simplices
In this section we introduce three quantities related to a simplex, which tell us about the size and shape of a simplex.
To start with, given a simplex, we are interested in the maximal ($\operatorname{{r_{\mathrm{max}}}}$) and minimal ($\operatorname{{r_{\mathrm{min}}}}$) distances between a vertex and the face opposite to it. The former quantity controls for instance how well a map is approximated by its linearization with respect to a given triangulation (see Section 4). Their ratio on the other hand tells us how degenerate the simplex is.
For $\operatorname{{r_{\mathrm{max}}}}$ we note that the maximum distance between a vertex and opposite face agrees with the maximum length of the edges adjacent to both the vertex and the opposite face. For $\operatorname{{r_{\mathrm{min}}}}$ we interpret the distance between vertex and opposite face as the distance between the vertex and the affine plane spanned by the face. Below we define $\operatorname{{r_{\mathrm{max}}}}$ and $\operatorname{{r_{\mathrm{min}}}}$ for $(m+1)$-tuples of points in ${\mathbb{R}}^{N}$, but we can also speak of $\operatorname{{r_{\mathrm{max}}}}$ and $\operatorname{{r_{\mathrm{min}}}}$ of a linear simplex $\Delta$ by identifying $\Delta$ with its vertices.
@@
The functions
$\operatorname{{r_{\mathrm{min}}}},\operatorname{{r_{\mathrm{max}}}}:{\mathbb{R}}^{N}\times\dots\times{\mathbb{R}}^{N}={\mathbb{R}}^{N(m+1)}\to{\mathbb{R}}$
are defined by
$\operatorname{{r_{\mathrm{min}}}}(p_{0},\dots,p_{m})$ $=\min_{i\in[m]}d\left(p_{i},\operatorname{ASpan}\left(\langle p_{0},\dots,\hat{p_{i}},\dots,p_{m}\rangle\right)\right)\text{ \quad and}$
$\operatorname{{r_{\mathrm{max}}}}(p_{0},\dots,p_{m})$ $=\max_{i,j\in[m]}d(p_{i},p_{j})$
We recall from Definition 2.16 that each $m$-simplex $\Delta$ in ${\mathbb{R}}^{N}$ defines a unique linear map $T_{\Delta}:{\mathbb{R}}^{m}\to{\mathbb{R}}^{N}$, along with its inverse $T_{\Delta}^{-1}$. Therefore, we can also study the shape of $\Delta$ by studying the operators $T_{\Delta}$ and $T_{\Delta}^{-1}$, and in particular their norm. We bound in Lemma 3.7 the norm of $T_{\Delta}$ using $\operatorname{{r_{\mathrm{max}}}}$, but for $\|T_{\Delta}^{-1}\|$ we introduce an additional quantity $\Lambda$}. As in the case of $\operatorname{{r_{\mathrm{max}}}}$ and $\operatorname{{r_{\mathrm{min}}}}$, we define $\Lambda$} for $(m+1)$-tuples of points in ${\mathbb{R}}^{N}$. The difference is however that here we need to assume that the simplex $\Delta$ is ordered to be able to speak of ${\Lambda}(\Delta)$, since $\Lambda$} is only invariant under permutations of its input fixing the first element.
Definition 3.6.
The function
$\Lambda:\left\{(v_{0},\dots,v_{m})\in{\mathbb{R}}^{N}\times\dots\times{\mathbb{R}}^{N}={\mathbb{R}}^{N(m+1)}\mid v_{i}\neq v_{j}\text{ for all }i\neq j\right\}\to{\mathbb{R}}$
is defined by
$\Lambda(v_{0},\dots,v_{m})=\max_{\begin{subarray}{c}\lambda_{1},\dots,\lambda_{m}\in{\mathbb{R}}\\ |\sum\lambda_{i}(v_{i}-v_{0})|=1\end{subarray}}|\lambda_{i}|.$
We point out that if $v_{0}=0$ and the points $v_{1},\dots,v_{m}$ form an orthogonal frame, the quantity $\Lambda(v_{0},\dots,v_{m})$ equals $\operatorname{{r_{\mathrm{min}}}}(v_{0},\dots,v_{m}).11$1For the interested reader, we note that a closed form expression can be obtained for $\Lambda$} by using the Lagrange multiplier method. If $m=2$ and $v_{0}=0$, we obtain for instance that $\Lambda(0,v_{1},v_{2})=\frac{\max\{|v_{1}|,|v_{2}|\}}{\sqrt{|v_{1}|^{2}|v_{2}|^{2}-\langle v_{1},v_{2}\rangle^{2}}}.$ In this case we see that the denominator resembles the Cauchy-Schwarz inequality and hence measures the degeneracy of the simplex $\langle 0,v_{1},v_{2}\rangle$.
Having defined the quantity $\Lambda$}, we obtain the following bounds on the linear maps associated to an ordered, linear simplex.
@@
Let $\Delta$ be an ordered, linear $m$-simplex in ${\mathbb{R}}^{N}$, and let $T_{\Delta}:{\mathbb{R}}^{m}\to{\mathbb{R}}^{N}$ be its associated linear map. Then the following bounds hold:
  • $\|T_{\Delta}\|\leq m\cdot\operatorname{{r_{\mathrm{max}}}}(\Delta)$, and
  • $\|T_{\Delta}^{-1}\|=m\cdot{\Lambda}(\Delta)$.
Proof.
We write the standard simplex $\Delta^{m}$ as $\langle e_{0},\dots,e_{m}\rangle$ and the simplex $\Delta$ as $\langle v_{0},\dots,v_{m}\rangle$. We assume without loss of generality that $0=e_{0}=v_{0}$. Then the map $T_{\Delta}$ is defined by $e_{i}\mapsto v_{i}$. We observe that
$\|T_{\Delta}\|=\max_{\begin{subarray}{c}\mu_{1}\dots,\mu_{m}\in{\mathbb{R}},\\ |\sum\mu_{i}e_{i}|=1\end{subarray}}\left|T_{\Delta}\left(\sum_{i=1}^{m}\mu_{i}e_{i}\right)\right|\leq\max_{\begin{subarray}{c}\mu_{1}\dots,\mu_{m}\in{\mathbb{R}},\\ |\sum\mu_{i}e_{i}|=1\end{subarray}}\sum_{i=1}^{m}|\mu_{i}||v_{i}|\leq\sum_{i=1}^{m}1\cdot\operatorname{{r_{\mathrm{max}}}}(\Delta)=m\cdot\operatorname{{r_{\mathrm{max}}}}(\Delta)$
and similarly that
$\|T_{\Delta}^{-1}\|$ $=\max_{\begin{subarray}{c}\mu_{1}\dots,\mu_{m}\in{\mathbb{R}},\\ |\sum\mu_{i}v_{i}|=1\end{subarray}}\left|T_{\Delta}^{-1}\left(\sum_{i=1}^{m}\mu_{i}v_{i}\right)\right|\leq\max_{\begin{subarray}{c}\mu_{1}\dots,\mu_{m}\in{\mathbb{R}},\\ |\sum\mu_{i}v_{i}|=1\end{subarray}}\sum_{i=1}^{m}|\mu_{i}||e_{i}|\leq\sum_{i=1}^{m}\max_{\begin{subarray}{c}\mu_{1}\dots,\mu_{m}\in{\mathbb{R}},\\ |\sum\mu_{i}v_{i}|=1\end{subarray}}|\mu_{i}|=m\cdot{\Lambda}(\Delta).\qed$
A priori we cannot bound the quantities $\operatorname{{r_{\mathrm{min}}}}$, $\operatorname{{r_{\mathrm{max}}}}$ and $\Lambda$} among all subdivisions of a linear polyhedron, since the space of all linear simplices in ${\mathbb{R}}^{N}$ is not compact. However, due to the existence of the model simplices (from Lemma 3.4), we can bound these quantities when considering crystalline subdivisions. We point out that the product of $\operatorname{{r_{\mathrm{max}}}}$ and $\Lambda$} is in particular independent of the number of subdivisions. Additionally, we observe that $\operatorname{{r_{\mathrm{min}}}}$ and $\operatorname{{r_{\mathrm{max}}}}$ decrease when we apply crystalline subdivisions, whereas $\Lambda$ increases.
Lemma 3.8.
Let $K$ be simplicial complex that is ordered and finite. Then, there exists $B,C,D,E\in{\mathbb{R}}_{+}$ such that for all $\ell\in\mathbb{N}$:
$\min_{\Delta\in K_{\ell}}\operatorname{{r_{\mathrm{min}}}}(\Delta)=B\cdot 2^{-\ell}$ $,\quad\max_{\Delta\in K_{\ell}}\operatorname{{r_{\mathrm{max}}}}(\Delta)=C\cdot 2^{-\ell}$
$\max_{\Delta\in K_{\ell}}\Lambda(\Delta)=D\cdot 2^{\ell}\quad\textrm{ and }$ $\quad\max_{\Delta\in K_{\ell}}\operatorname{{r_{\mathrm{max}}}}(\Delta)\cdot{\Lambda}(\Delta)=CD.$